OpenStudy (anonymous):

You just purchased a cellular phone and are trying to determine to which cell phone company you will give your business. When you contacted the Talks-A-Lot Company, they were offering a monthly plan of $40 for 600 minutes and $0.35 for each minute exceeding the 600 minutes. In the Sunday paper you see an ad for the Chat-Away Company, which offers a monthly plan of $50 for 600 minutes and $0.10 for each minute exceeding the 600 minutes. How many minutes would you have to talk over and above the 600 minutes for the cost to be the same with both companies? What would be the equal cost?

OpenStudy (anonymous):

let C = cost m = minutes for talk a lot we have a point (40, 600) and it has to increase 0.35 by minute so its slope is 0.35, using the equation for the line we have C-40 =0.35(x-600)

OpenStudy (anonymous):

c-40=.35x-210

OpenStudy (anonymous):

doint the same with the chat away company we have (C-50) = 0.1 (x-600)

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

u have to find the intersection point of both lines

OpenStudy (anonymous):

c-40=.35x-210 c-50=.1x-60

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

find the intersection point

OpenStudy (anonymous):

what do i do next

OpenStudy (anonymous):

well u could symplify this equation c-40=.35x-210 like c=0.35x -170 and the another like c= 0.1x -10

OpenStudy (anonymous):

now make equal both equations 0.35x -170 =0.1x -10 and solve for x

OpenStudy (anonymous):

x=640